OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. (2,-1);2x-9y=8

OpenStudy (mimi_x3):

m = 2/9 y - 1 = 2/9(x-2) 9y-9 = 2x - 4 2x - 9y + 5 = 0

OpenStudy (anonymous):

y=(2/9)x+13/9

OpenStudy (ybarrap):

2x - 9y=8 -9y = 8 - 2x 9y = 2x - 8 y = 2x/9 - 8/9 slope = 2/9 y-intercept = -8/9 ynew = slope*x + b2 ynew(2) = -1 = 2(2)/9 - b2, where b2 is y-intercept of parallel line going through point (2,-1) b2 = 4/9 + 1 = (4 + 9)/9 = 13/9 Ans. ynew = 2x/9 + 13/9

OpenStudy (anonymous):

the slope is m=2/9 then sub x=2, y=-1 and m=2/9 into y=mx+b to get b -1=(2/9)(2)+b -1=4/9 + b b=-13/9 now just sub b=-13/9 and m=2/9 into y=mx+b y=(2/9)x-13/9

OpenStudy (ybarrap):

Note b2 is -13/9 not 13/9.

OpenStudy (anonymous):

thank you all for your help

OpenStudy (mimi_x3):

woops my answer is wrong then xD

OpenStudy (anonymous):

@mimi you made a sign error on the LHS of point-slope: y-(-1)

OpenStudy (mimi_x3):

oh yeah woops my bad, so~ sleepy xD